[Brahms-l] Shift report 20040117 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 17 2004 - 07:42:37 EST
------------------------------------------------------------
Date: Sat Jan 17 07:42:36 2004, Shift: 00:00-08:00
Supervisor: Ionut ARSENE, with: Sorin ZGURA
Log book#15 pages 34 - 34

Shift Summary:
------------------------------------------------------------
1. Beam for almost the entire night shift.
   Problems with D2 (water leakage). D2 was stopped (see E-log)

2. Settings: FS@12deg. with D1@B430 & MRS@90deg. with D5@A2255
   Data runs:
  9719 - 150k trig6 & 47k trig3 
  9720 - 100k trig6 & 32k trig3
  9721 - 100k trig6 & 70k trig3
  9723 - 0k trig6   & 50k trig3
  9724 - still running

3. FS was taken out for runs 9723 and 9724.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 17 07:42:57 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 17 2004 - 07:43:16 EST